# Prospect Theory Fails for LLMs: Revealing Instability of Decision-Making under Epistemic Uncertainty

### v1

@@ Appendix A Lottery Design
Table 5, 6 and 7 shows the lottery design for PT parameter estimation. The values here are from Tanaka et al. (2010). They are specially designed to get best PT parameters.
@@
Appendix B Prompt Design
Figure 7 shows prompt design for decision-making evaluation test and probability mapping test. All lotteries are sampled 256 times. To simulate human decision-making, while keeping the model directly output its answer, up to 15 history decisions are maintained. Meanwhile, we use a random order for the 35 lotteries to relieve positional bias. An introduction is provided at the very beginning. NUMBER will be replaced by the values stated in table 5, 6 and 7 .
@@ Appendix B Prompt Design
Appendix C Detailed Experimental Results
@@
Specifically, Tables 8, 9, and 10 report the estimates of parameters $\sigma$, $\lambda$, and $\gamma$ with their 95% confidence intervals for each model and round. Table 11 summarizes the models’ mean absolute errors (MAE) and McFadden $R^{2}$ values across rounds.
@@
Appendix D Discussion and Implications
@@
Appendix E Failure Cases Analysis
@@
We summarize two major issues observed in the experimental outcomes:
Figure 9: An example of non-monotonic result with multiple switching points. This result comes from marker “It’s undoubtedly to” with Qwen2.5-7B-Instruct model over 256 trials. The blue line shows counts of option K selections, and the yellow dashed line marks half the trials. The crosses denote switching points where choices for K and U are equal.
@@
(2) Shift from first-person to third-person framing. The original markers in Zhou et al. (2024) were presented in the first person (e.g., “I am not confident, maybe it’s…”), whereas our experiment reformulated them into third-person expressions (e.g., “It’s not confident, maybe can…”). This change in perspective may alter the perceived epistemic force of the marker for the model.
(3) Intrinsic instability of epistemic markers. Even for human interpretation, such markers are context-dependent and inherently imprecise Liu et al. (2025b); Guo et al. (2025); Wang et al. (2025); Zheng et al. (2025a, b). Their probability mapping by LLMs in economic decision-making contexts may exhibit fundamental reliability flaws Zheng et al. (2025c).
These limitations motivated the redesign of our marker set and prompt formulation in subsequent experiments, aiming for shorter, syntactically simpler, and context-consistent expressions.

### v2

@@ Appendix B Lottery Design
Table 7, 8 and 9 shows the lottery design for PT parameter estimation. The values here are from Tanaka et al. (2010). They are specially designed to get best PT parameters.
Appendix C Prompt Design
Figure 7 shows prompt design for decision-making evaluation test and probability mapping test. All lotteries are sampled 256 times. To simulate human decision-making, while keeping the model directly output its answer, up to 15 history decisions are maintained. Meanwhile, we use a random order for the 35 lotteries to relieve positional bias. An introduction is provided at the very beginning. NUMBER will be replaced by the values stated in table 7, 8 and 9 .
Appendix D Marker Replacement Rules
For the details of how we replace probabilities with markers, see Table 6.
Model 30% 70% 10% 90%
Qwen2.5-7B-Instruct uncertain almost certain somewhat likely highly likely
Llama3.1-8B-Instruct likely almost certain very unlikely almost certain
Mistral-7B-Instruct-v0.3 very unlikely highly likely highly unlikely almost certain
Qwen2.5-14B-Instruct somewhat unlikely highly likely very unlikely almost certain
Qwen2.5-32B-Instruct somewhat unlikely probable somewhat likely almost certain
Table 6: Marker Replacement Rules for Different Models. This is determined for introducing the least numeric differences, and balancing model and human interpretations.
@@ Appendix D Marker Replacement Rules
Appendix E Detailed Experimental Results
@@
Specifically, Tables 10, 11, and 12 report the estimates of parameters $\sigma$, $\lambda$, and $\gamma$ with their 95% confidence intervals for each model and round. Table 13 summarizes the models’ mean absolute errors (MAE) and McFadden $R^{2}$ values across rounds.
@@
Appendix F Discussion and Implications
@@
Appendix G Failure Cases Analysis
@@
We summarize two major issues observed in the experimental outcomes:
Figure 8: An example of non-monotonic result with multiple switching points. This result comes from marker “It’s undoubtedly to” with Qwen2.5-7B-Instruct model over 256 trials. The blue line shows counts of option K selections, and the yellow dashed line marks half the trials. The crosses denote switching points.
@@
(2) Shift from first-person to third-person framing. The original markers in Zhou et al. (2024) were presented in the first person (e.g., “I am not confident, maybe it’s…”), whereas our experiment reformulated them into third-person expressions (e.g., “It’s not confident, maybe can…”).
(3) Intrinsic instability of epistemic markers. Even for human interpretation, such markers are context-dependent and inherently imprecise Liu et al. (2025c); Wang et al. (2025b); Zheng et al. (2025a, b); Zong et al. (2025a); Guo et al. (2025). Their probability mapping by LLMs in economic decision-making contexts may exhibit fundamental reliability flaws Zheng et al. (2025c); Liu et al. (2024).
These limitations motivated the redesign of our marker set and prompt formulation in subsequent experiments.